8. A solution is prepared by adding 0.10 mol oflithium nitrate, LiNO, to 1.00 Lof water. Which statement about the solution is correct? A)The solution is basic. B) The solution is C) The solution is weakly acidic. D) The solution is strongly acidic. E The values for K and K, for the species in solution must be known before a prediction can be made. The pH of blood is 7.35. It is maintained in part by the buffer system composed of carbonic acid H CO) and the bicarbonate (hydrogen carbonate, HCO, ion. What is the ratio of bicarbonateyIcarbonic acid] at this pH? For carbonic acid, K. 4.2 x 10. A) (bicarbonate)llcarbonic acid] 0.11 D) Ubicarbonatey[carbonic acid 9.4 B) (bicarbonate)llcarbonic acid)-0.38 E) None of the above ratios is correct. C bicarbonate Mcarbonic acid] 2.65 10. A diprotic acid H2A has Kal 1 10 and Ke 1 0 The corresponding base A is titrated with aqueous HCI, both solutions being 0.1 molL .Which one of the following diagrams best represents the titration curve which will be seen? Truel False (2pts each) 1. Increasing the concentrations of the components of a buffer solution will increase the buffer range. 2. the pH of a buffer solution is greater than the pk.value of the buffer acid, the buffer will have more capacity to neutralize added base than added acid. 3.The equivalence point in a titration is defined as the point when the indicator changes color. 4. For a diprotic acid HA. the relationship Ku K is always true.
Answer #1

(8) Since LiNO3 is a salt formed by stron base (LiOH) and strong acid (HNO3), when it dissolved in water it forms a neutral solution because both can undergo complete ionisation. No hydrolysis takes place.

(B)

(9)

pH = pKa + Lg[salt]/[acid]

7.35 = - Log(4.2*10-7) + Log[salt]/[acid]

7.35 = 6.38 + Log[salt]/[acid]

Log[salt]/[acid] = 0.970

[salt]/[acid] = 9.33

(D)